OpenStudy (anonymous):

2m-2x=x+n. solving for x

OpenStudy (texaschic101):

2m - 2x = x + n (add 2x to both sides) 2m - 2x + 2x = 2x + x + n 2m = 3x + n (now subtract n from both sides) 2m - n = 3x (now divide both sides by 3) (2m - n)/3 = x or 2/3m - 1/3n = x

OpenStudy (anonymous):

Thank you for showing me step by step how to solve this. very helpful.